11.
In a bell crank lever reaction at fulcrum is 12 kN , if permissible bearing stress is 8MPa calculate diameter of pin.
1
Expert's answer
2021-09-29T04:38:29-0400

"F = A × P\\
F = π ×(d²/4) × P\\"


"d =\sqrt{\dfrac{ 4F}{π×P }}=\sqrt\dfrac{4×12×10³}{π×8×10⁶} ="

"0.044m = 4.4cm"